# The Arens–Michael envelope of a solvable Lie algebra is a homological epimorphism

### v1

@@ Introduction
We are interested in the question under what conditions, for a given finite-dimensional complex Lie algebra $\mathfrak{g}$, the universal completion (called the Arens–Michael envelope) of $U(\mathfrak{g})$ (its universal enveloping algebra) is a homological epimorphism. This question is implicitly contained in the works of J. Taylor published in early 70s [32, 33, 34, 35] and explicitly formulated by Pirkovskii in [27, § 9, Problem 1].
@@
Flat homomorphisms play an important role in algebra and analysis but in some respects the requirement of flatness is too strong and must be replaced by a weaker condition. The concept of homological epimorphism was introduced J. Taylor [34] as a weakened form of flatness under the name of “absolute localization” and has been rediscovered several times in different contexts and under different names — “lifting”, “stably flat homomorphism”, “isocohomological morphism”, “homotopy epimorphism”; see references in [10, Remark 3.16]. An important property for applications is that homological epimorphisms preserve homologies and cohomologies [34, Propositions 1.4 and 1.7]. We are interested in the functional analytic version of this concept that comes back to Taylor and consider homological epimorphisms of complete locally convex algebras with jointly continuous multiplication (we call them $\mathop{\widehat{\otimes}}$-algebras); see Definition 4.2.
An example of the use of homological epimorphisms is contained in Meyer’s preprint [24], where Connes’s calculation of cyclic cohomology of a smooth non-commutative torus in [15, § 6] is simplified. In addition, it was noted by Pirkovskii in [25] that flat homomorphisms are not sufficient for characterization of open embeddings of Stein varieties but a weak version of the notion of homological epimorphism is adequate for this task; for an extension of results to general Stein spaces see [10, 13]. Another motivation for studying of homological epimorphisms in the context of functional analysis comes from non-commutative spectral theory; see the introduction to [27] for a detailed discussion and interpretation of Taylor’s ideas.
@@
The idea to study Arens–Michael envelopes and, in particular, to look for conditions under which they are homological homomorphisms also belongs to Taylor. Recall that the Arens–Michael envelope of a topological algebra $A$ is the universal completion with respect to the class of Banach algebras (equivalently, the completion relative to the locally convex topology determined by all continuous submultiplicative seminorms on $A$). Roughly speaking, this completion is responsible for continuous representations of $A$ on Banach spaces; see details in § 2. A prototype example of an Arens–Michael envelope is the embedding of the polynomial algebra $\mathbb{C}[z_{1},\ldots,z_{n}]$ to the algebra of entire functions $\mathcal{O}(\mathbb{C}^{n})$.
@@
Let $U(\mathfrak{g})$ be the universal enveloping algebra of a finite-dimensional complex Lie algebra $\mathfrak{g}$. The question of whether the Arens–Michael envelope homomorphism $U(\mathfrak{g})\to\widehat{U}(\mathfrak{g})$ is a homological epimorphism has been studied since 70s. The affirmarive answer in the case when $\mathfrak{g}$ is abelian, and the negative answer in the case when it is semisimple, was given by Taylor in [34]. Three decades later, in [26], Pirkovskii showed that the solvability a necessary condition. At the same time, the study of the nilpotent case was started by Dosiev [16, 17]. It was continued by Pirkovskii in [27] and recently completed by the author in [2], where it was shown that the nilpotency is a sufficient condition. Thus, only the case when $\mathfrak{g}$ is solvable but not nilpotent remained uninvestigated (with one exception — the two-dimensional non-abelian algebra; see [27]). Here we prove that the solvability is also a sufficient condition; see Theorem 4.4. Thus, the final answer is that $U(\mathfrak{g})\to\widehat{U}(\mathfrak{g})$ is a homological epimorphism if and only if $\mathfrak{g}$ is solvable.
@@
  • In [2], the nilpotent case is considered. The structure of $\widehat{U}(\mathfrak{g})$ is described and proved that Arens–Michael envelope is a homological epimorphism (using a method different from that applied in this article).
  • In [3], the structure of the algebra of analytic functionals on a connected complex Lie group is discussed. In particular, the results there can be applied to $\widehat{U}(\mathfrak{g})$ for a general finite-dimensional Lie algebra. Some improvements is obtained in [8].
  • A decomposition into iterated analytic smash products (which is essential here) is constructed in [9].
  • The decomposition mentioned above exists not only for the Arens–Michael envelope but for other completions. Some preliminary work for these completions is done in [6, 7]. Formally, [9] is based on those preliminary results but, in fact, the case of the Arens–Michael envelope can be studied without reference to [6, 7].
Our main result is of independent interest but, on the other hand, this text is a part of a big project on completions of universal enveloping algebras and algebras of analytic functionals. Homological properties related with the first topic (universal enveloping algebras) is a subject of this paper. The author plans to study homological properties of algebras of analytic functionals and their completions in a separate article.
The result was announced at the seminar “Algebras in Analysis” in Moscow State University in 2018. The author later discovered a gap in the proof proposed at the time. This text contains an argument in which the gap is filled.
@@ 1. Main ideas of the proof
The proof is based on two circles of ideas: author’s results on decomposition of $\widehat{U}(\mathfrak{g})$ into an iterated analytic smash products obtained in [9] and a powerful technique of relative homological epimorphisms developed by Pirkovskii in [28]. In fact, the analytic part of preliminary work is contained in [9] and preceding articles. Pirkovskii’s technique, which is applied below, is more algebraic in nature.
When studying homological epimorphisms, the question of finding a convenient projective resolution is important. To construct such a resolution one can use a decomposition of $U(\mathfrak{g})$ into an iterated Ore extension. In group homology, a construction of a resolution of the trivial module of a semi-direct product is well known (see, e.g., [21, Chapter 5, § 2, pp. 250–251]). It is not hard to generalize this construction for cocommutative Hopf algebras and transfer it into the context of functional analysis. (This topic will be discussed elsewhere.) Unfortunately, it is not clear whether or not the tensor product functor (4.1) sends such projective resolution to a similar resolution. In [28], to overcome this kind of difficulty it was proposed to use relative homological homomorphisms instead of usual. We show below that, with an appropriate modification, this idea works well in our case.
@@
In [28, Theorem 9.12], Pirkovskii proved that the Arens–Michael envelope is a homological epimorphism for a number of finitely-generated non-commutative algebras. In this paper, we are interested only in generalizing of Parts (vi) and (vii) of the above theorem, the parts that concern the two-dimensional non-abelian and three-dimensional Heisenberg Lie algebra. This section contains an extract of argument in these cases. We postpone the necessary definitions until § 4.
In both the two-dimensional and three-dimensional cases considered in [28], a decomposition of $\mathfrak{g}$ into a semidirect sum induces a decomposition of $U(\mathfrak{g})$ into an Ore extension of a commutative algebra. Namely, $U(\mathfrak{g})$ is isomorphic to $\mathbb{C}[x][t;\delta]$, where $\delta=y\frac{d}{dy}$, in the first case and $\mathbb{C}[x,y][t;\delta]$, where $\delta=y\frac{\partial}{\partial x}$ in the second. As a corollary, $\widehat{U}(\mathfrak{g})$ can be represented as an analytic Ore extension $\mathcal{O}(\mathbb{C},R_{\delta};\widehat{\delta})$, where $R_{\delta}$ is a certain completion of $R$ and $\widehat{\delta}$ is the extension $\delta$ to $R_{\delta}$ [28, Theorem 5.1]. (Here and below $\mathcal{O}$ stands for holomorphic functions.)
Pirkovskii used a two-step argument. First he showed that $R\to R_{\delta}$ is a homological homomorphism and next that $\iota$ is a one-sided relative homological homomorphism with respect to $R_{\delta}$. Both conditions are included into the hypotheses of the following result.
@@
[28, Theorem 9.1(ii)] Let $(f,g)\!:(A,R)\to(B,S)$ be an $R$-$S$-homomorphism from an $R$$\mathop{\widehat{\otimes}}$-algebra $A$ to an $S$$\mathop{\widehat{\otimes}}$-algebra $B$. Suppose that
@@
In the case of low-dimensional Lie algebras, consider $R$-$R_{\delta}$-homomorphism
@@
where $g$ denotes the $R\to R_{\delta}$. To prove that $\iota$ is a homological epimorphism it suffices to check Conditions (1)–(4) in Theorem 1.1 for $(\iota,g)$.
For Ore extensions and analytical Ore extensions, Condition (3) easily follows from definitions. An Ore extension of $R$ is relatively $(f_{2})$-quasi-free $R$-algebra [28, Example 8.1], which in particular implies that it is of $(f_{2})$-finite type [28, Definition 8.5 and Proposition 7.3.]. Thus Condition (4) holds. (Note that the use of relatively quasi-free algebras was an important innovation in this topic.) Checking of Conditions (1) and (2) is more challenging.
To prove Condition (2) Pirkovskii applied the following result.
@@
[28, Theorem 7.6(ii)] Let $A$, $R$, $S$ be $\mathop{\widehat{\otimes}}$-algebras and $g\!:R\to S$ an epimorphism of $\mathop{\widehat{\otimes}}$-algebras. Suppose that $A$ is an $R$$\mathop{\widehat{\otimes}}$-algebra and its Arens–Michael envelope $\widehat{A}$ is an $S$$\mathop{\widehat{\otimes}}$-algebra such that the pair $(\iota_{A},g)$ is an $R$-$S$-homomorphism from $A$ to $\widehat{A}$. Assume also that $A$ is relatively quasi-free over $R$. Then $\iota_{A}\!:A\to\widehat{A}$ is a two-sided relative homological epimorphism.
It is essential in the proof that $\iota_{A}\!:A\to\widehat{A}$ satisfies to the unique extension property for derivations (Property (UDE)); see the corresponding definition in § 3.
@@
In the high dimensional case, it is natural to use induction. Specifically, if an algebra $A$ has a chain of subalgebras $\mathbb{C}=R_{0}\subset R_{1}\subset\cdots\subset R_{n}=A$, where each of $R_{i}$ is relatively quasi-free over $R_{i-1}$, then an iterative application of Theorem 1.1 enable us with a tool for proving that $A\to\widehat{A}$ is a homological epimorphism; see the introduction to [28]. This plan works well in some situations but unfortunately cannot be applied to all high-dimensional solvable Lie algebras. Indeed, a finite-dimensional solvable complex Lie algebra $\mathfrak{g}$ admits an iterated semidirect sum decomposition,
@@
where $\mathfrak{f}_{1},\ldots,\mathfrak{f}_{n}$ are $1$-dimensional. Moreover, this decomposition actually induces an decomposition of the universal enveloping algebra into an iterated Ore extension; see, e.g., [23, pp. 33–34, 1.7.11(iv)]. Nevertheless, in this case, $R_{n-1}\to(R_{n-1})_{\delta}$ is not usually an Arens–Michael envelope and, moreover, $(R_{n-1})_{\delta}$ cannot be represented as an analytic Ore extension.
@@
Our strategy is to apply Theorem 1.1 using iterations. But the scheme proposed in [28] needs to be modified, as is obvious from previous considerations.
1. We use iterated analytic smash products instead of analytic Ore extensions; see Theorem 2.5. An analytic Ore extension with trivial twisting is an analytic smash product but when the twisting is non-trivial this is not the case. (Thus most of parts of Theorem 9.12 in [28] are not covered by our approach.) Some preliminaries on Hopf $\mathop{\widehat{\otimes}}$-algebras, analytic smash products and Arens–Michael envelopes is contained in § 2.
2. We use completions of $\mathbb{C}[z]$ other than $\mathcal{O}(\mathbb{C})$, the algebras $\mathfrak{A}_{s}$; see definition in (2.7). The essential point is that $\mathcal{O}(\mathbb{C})\to\mathfrak{A}_{s}$ and some induced homomorphisms of analytic Ore extensions satisfy Property (UDE). To prove theses facts we apply an auxiliary result on pushouts; see Theorem 3.4 and preliminary results in § 2.
3. We get an Arens–Michael envelope only on the final step of iteration. So we have to replace Theorem 1.2 by a more general result on homomorphisms satisfying Property (UDE), Theorem 4.7. Combining it with a decomposition result essentially proved in [9] (Theorem 2.5), we finally deduce Theorem 4.4.
@@ 2. Hopf $\mathop{\widehat{\otimes}}$-algebras, analytic smash products and Arens–Michael envelopes
Consider the bifunctor $(-)\mathbin{\widehat{\otimes}}(-)$ of complete projective tensor product on the category of complete locally convex spaces and Hopf algebras in the corresponding symmetric monoidal category. We call them Hopf $\mathbin{\widehat{\otimes}}$-algebras (read ‘topological Hopf algebras’); see [22] or [27]. Also, $\mathbin{\widehat{\otimes}}$-algebras and $\mathbin{\widehat{\otimes}}$-(bi)modules, i.e., complete locally convex algebras and (bi)modules with jointly continuous multiplication, are considered. We assume that each algebra contains an identity and each module is unital. A $\mathbin{\widehat{\otimes}}$-(bi)module over a $\mathbin{\widehat{\otimes}}$-algebra $A$ is referred as an $A$$\mathbin{\widehat{\otimes}}$-(bi)module. We also assume that $\mathop{\widehat{\otimes}}$-algebra homomorphisms preserve identity and are continuous (as well as $\mathbin{\widehat{\otimes}}$-module morphisms).
@@
The Sweedler notation is widely used in the Hopf algebra theory instead of the tensor notation because the latter is not always convenient. It was noted in [1, § 2.4] that this notation can be generalized to topological Hopf algebras. A version of the generalized Sweedler notation sufficient for our purposes is described in [9] and we briefly recall it here.
In the classical Sweedler notation, the comultiplication $\Delta$ on a Hopf algebra has the form
$\Delta(h)=\sum h_{(1)}\otimes h_{(2)}.$ (2.1)
Here an arbitrary representation of $\Delta(h)$ is taken.
@@
In the case of a Hopf $\mathop{\widehat{\otimes}}$-algebra, (2.1) is a short notation for such a representation.
@@
In the generalized Sweedler notation, it takes the same form as in the classical Sweedler notation:
@@
By the iterated limit theorem [20, Chapter 2, p. 69, Theorem 4], both iterated limits in (2.2) can be replaced by a simple limit of a net. In particular, this means that the iteration of the comultipication can be written as
@@
Also, in this notation, the antipode axiom for a Hopf $\mathop{\widehat{\otimes}}$-algebra takes the form
@@
Let $H$ be a Hopf $\mathbin{\widehat{\otimes}}$-algebra and $A$ a $\mathbin{\widehat{\otimes}}$-algebra that is a left $H$$\mathbin{\widehat{\otimes}}$-module. Then $A\mathbin{\widehat{\otimes}}A$ and $\mathbb{C}$ are left $H$$\mathbin{\widehat{\otimes}}$-modules with respect to
@@
For a $\mathbin{\widehat{\otimes}}$-algebra homomorphism $\psi\!:H\to B$ consider the adjoint action $H$ on $B$:
@@
Let $H$ be a Hopf $\mathbin{\widehat{\otimes}}$-algebra and $A$ an $H$$\mathbin{\widehat{\otimes}}$-module algebra. The analytic smash product $A\mathop{\widehat{\#}}H$ is defined as a $\mathbin{\widehat{\otimes}}$-algebra endowed with $\mathbin{\widehat{\otimes}}$-algebra homomorphisms $i\!:A\to A\mathop{\widehat{\#}}H$ and $j\!:H\to A\mathop{\widehat{\#}}H$ such that the following conditions holds.
(A) $i$ is a $\mathbin{\widehat{\otimes}}$-module algebra homomorphisms with respect to the adjoint action associated with $j$.
@@
In the case when $H=\mathcal{O}(\mathbb{C})$ and act on $A$ by a derivation, the analytic smash product coincides with the corresponding analytic Ore extension [28, Remarks 4.4 and 4.2]. The same holds in the pure algebraic case, i.e., when $\mathbb{C}[z]$ acts by a derivation.
@@
$\mathfrak{A}_{s}\!:=\Bigl{\{}a=\sum_{n=0}^{\infty}a_{n}z^{n}\!:\|a\|_{r,s}\!:=\sum_{n=0}^{\infty}|a_{n}|\,\frac{r^{n}}{n!^{s}}<\infty\;\forall r>0\Bigr{\}},$ (2.7)
where $z$ is treated as a formal variable, and endow $\mathfrak{A}_{s}$ with the topology determined by the seminorm family $(\|\cdot\|_{r,s};\,r>0)$. Denote also $\mathbb{C}[[z]]$, the algebra of all power series in $z$, by $\mathfrak{A}_{\infty}$.

### v4

@@ Introduction
We are interested in the question under what conditions, for a given finite-dimensional complex Lie algebra $\mathfrak{g}$, the universal completion (called the Arens–Michael envelope) of $U(\mathfrak{g})$ (its universal enveloping algebra) is a homological epimorphism. This question is implicitly contained in a paper of J. Taylor published in early 70s [31] and explicitly formulated by Pirkovskii in [25, § 9, Problem 1].
@@
Flat homomorphisms play an important role in algebra and analysis but in some respects the requirement of flatness is too strong and must be replaced by a weaker condition. The concept of homological epimorphism was introduced by J. Taylor [31] as a weakened form of flatness under the name of “absolute localization” and has been rediscovered several times in different contexts and under different names — “lifting”, “stably flat homomorphism”, “isocohomological morphism”, “homotopy epimorphism”; see references in [10, Remark 3.16]. An important property for applications is that homological epimorphisms preserve Hochschild homologies and cohomologies [31, Propositions 1.4 and 1.7]. We are interested in the functional analytic version of this concept that goes back to Taylor and consider homological epimorphisms of complete locally convex algebras with jointly continuous multiplication (we call them $\mathop{\widehat{\otimes}}$-algebras); see Definition 4.2.
An example of the use of homological epimorphisms is contained in Meyer’s preprint [22], where Connes’s calculation of the cyclic cohomologies of smooth non-commutative tori in [13, § 6] is simplified. In addition, it was noted by Pirkovskii in [23] that flat homomorphisms are not sufficient for characterization of open embeddings of Stein varieties but a weak version of the notion of homological epimorphism is adequate for this task; for an extension of results to general Stein spaces see [10, 11]. Another motivation for studying of homological epimorphisms in the context of functional analysis comes from non-commutative spectral theory; see the introduction to [25] for a detailed discussion and interpretation of Taylor’s ideas.
@@
The idea to study Arens–Michael envelopes and, in particular, to look for conditions under which they are homological epimorphisms also belongs to Taylor. Recall that the Arens–Michael envelope of a topological algebra $A$ is the universal completion with respect to the class of Banach algebras (equivalently, the completion relative to the locally convex topology determined by all continuous submultiplicative seminorms on $A$); see details in § 2. Roughly speaking, this completion is responsible for continuous representations of $A$ on Banach spaces. A prototype example of an Arens–Michael envelope is the embedding of the polynomial algebra $\mathbb{C}[z_{1},\ldots,z_{n}]$ to the algebra of entire functions $\mathcal{O}(\mathbb{C}^{n})$.
@@
Let $U(\mathfrak{g})$ be the universal enveloping algebra of a finite-dimensional complex Lie algebra $\mathfrak{g}$. The question of whether the Arens–Michael envelope homomorphism $U(\mathfrak{g})\to\widehat{U}(\mathfrak{g})$ is a homological epimorphism has been studied since 70s. The affirmarive answer in the case when $\mathfrak{g}$ is abelian, and the negative answer in the case when it is semisimple, was given by Taylor in [31]. Three decades later, in [24], Pirkovskii showed that the solvability a necessary condition. At the same time, the study of the nilpotent case was started by Dosiev [14, 15]. It was continued by Pirkovskii in [25] and recently completed by the author in [2], where it was shown that the nilpotency is a sufficient condition. Thus, only the case when $\mathfrak{g}$ is solvable but not nilpotent remained uninvestigated (with one exception — the two-dimensional non-abelian algebra [25]). Here we prove that the solvability is also a sufficient condition; see Theorem 4.3. Thus, the final answer is that $U(\mathfrak{g})\to\widehat{U}(\mathfrak{g})$ is a homological epimorphism if and only if $\mathfrak{g}$ is solvable.
@@
  • In [2], the nilpotent case is considered. The structure of $\widehat{U}(\mathfrak{g})$ is described and proved that the Arens–Michael envelope is a homological epimorphism (using a method different from that applied in this article).
  • In [3], the structure of the algebra of analytic functionals on a connected complex Lie group is discussed. In particular, the results in [3] can be applied to $\widehat{U}(\mathfrak{g})$ for a general finite-dimensional Lie algebra. Some improvements is obtained in [8].
  • A decomposition into an iterated analytic smash products (which is essential here) is constructed in [9].
  • The decomposition mentioned above exists not only for the Arens–Michael envelope but for other completions. Some preliminary results on these completions is proved in [6, 7]. Formally, [9] is based on them but, in fact, the case of the Arens–Michael envelope can be studied without reference to [6, 7].
Our main result is of independent interest but, on the other hand, this text is a part of a big project on completions of universal enveloping algebras and algebras of analytic functionals. Homological properties related with the first topic is a subject of this paper. The author plans to study homological properties of algebras of analytic functionals and their completions in a separate article.
The result was announced at the seminar “Algebras in Analysis” at Moscow State University in 2018. The author later discovered a gap in the proof proposed at that time. This text presents a different argument.
@@ 1. Main ideas of the proof
The proof is based on two circles of sets: the author’s results on decomposing of $\widehat{U}(\mathfrak{g})$ into an iterated analytic smash product obtained in [9] and a powerful technique of relative homological epimorphisms developed by Pirkovskii in [26]. In fact, the analytic part of preliminary work is contained in [9] and preceding articles. The technique of Pirkovskii, which is applied below, is more algebraic in nature.
When studying homological epimorphisms, the question of finding a convenient projective resolution is important. To construct such a resolution one can use a decomposition of $U(\mathfrak{g})$ into an iterated smash product. In group homology, a construction of a resolution of the trivial module of a semi-direct product is well known (see, e.g., [19, Chapter 5, § 2, pp. 250–251]). It is not hard to generalize this construction to cocommutative Hopf algebras and transfer it to the context of functional analysis. (This topic will be discussed elsewhere.) Unfortunately, it is not clear whether or not the tensor product functor in (4.1) sends such projective resolution to a similar resolution. In [26], to overcome this kind of difficulties, it was proposed to use relative homological homomorphisms instead of usual. Below we show that, with appropriate modification, this idea works well in our case.
@@
In [26, Theorem 9.12], Pirkovskii proved that the Arens–Michael envelope is a homological epimorphism for a number of finitely-generated non-commutative algebras. In this paper, we are interested only in generalizing of Parts (vi) and (vii) of the above theorem, the parts that concern the two-dimensional non-abelian and three-dimensional Heisenberg Lie algebra. This section contains an extract of the argument in these cases. We postpone the necessary definitions until § 4.
In both the two-dimensional and three-dimensional cases considered in [26], a decomposition of $\mathfrak{g}$ into a semidirect sum induces a decomposition of $U(\mathfrak{g})$ into an Ore extension of a commutative algebra $R$. Namely, $U(\mathfrak{g})$ is isomorphic to $\mathbb{C}[y][x;\delta]$ with $\delta=y\frac{d}{dy}$ in the first case and $\mathbb{C}[y,z][x;\delta]$ with $\delta=z\frac{\partial}{\partial y}$ in the second. As a corollary, $\widehat{U}(\mathfrak{g})$ can be represented as an analytic Ore extension $\mathcal{O}(\mathbb{C},R_{\delta};\widehat{\delta})$, where $R_{\delta}$ is a certain completion of $R$ and $\widehat{\delta}$ is the extension of $\delta$ to $R_{\delta}$ [26, Theorem 5.1]. (Here and below $\mathcal{O}$ stands for holomorphic functions.)
Pirkovskii used a two-step argument. First he showed that $R\to R_{\delta}$ is a homological epimorphism and next that the Arens–Michael envelope homomorphism $\iota\!:U(\mathfrak{g})\to\widehat{U}(\mathfrak{g})$ is a one-sided relative homological epimorphism with respect to $R_{\delta}$. Both conditions are included into the hypotheses of the following result.
@@
[26, Theorem 9.1(ii)] Let $(f,g)\!:(A,R)\to(B,S)$ be an $R$-$S$-homomorphism from an $R$$\mathop{\widehat{\otimes}}$-algebra $A$ to an $S$$\mathop{\widehat{\otimes}}$-algebra $B$. Suppose that
@@
Returning to low-dimensional Lie algebras, consider $R$-$R_{\delta}$-homomorphism
@@
where $g$ denotes the homomorphism $R\to R_{\delta}$. To prove that $\iota$ is a homological epimorphism it suffices to check Conditions (1)–(4) in Theorem 1.1 for $(\iota,g)$.
For Ore extensions and analytical Ore extensions, Condition (3) easily follows from definitions. Note that Ore extensions have two useful properties: they are not only of $(f_{2})$-finite type but also relatively quasi-free; see Proposition 4.10 below. In particular, Condition (4) holds. Checking of Conditions (1) and (2) is more challenging. To prove Condition (2) Pirkovskii used relatively quasi-free algebras. (Note that the use of the relative quasi-freeness was a significant innovation in this topic.) In particular, the following result is important.
@@
[26, Theorem 7.6(ii)] Let $A$, $R$, $S$ be $\mathop{\widehat{\otimes}}$-algebras and $g\!:R\to S$ an epimorphism of $\mathop{\widehat{\otimes}}$-algebras and let $\iota_{A}\!:A\to\widehat{A}$ denote the Arens–Michael envelope. Suppose that $A$ is an $R$$\mathop{\widehat{\otimes}}$-algebra and $\widehat{A}$ is an $S$$\mathop{\widehat{\otimes}}$-algebra such that the pair $(\iota_{A},g)$ is an $R$-$S$-homomorphism. Assume also that $A$ is relatively quasi-free over $R$. Then $\iota_{A}$ is a two-sided relative homological epimorphism.
In the proof, it is essential that $\iota_{A}\!:A\to\widehat{A}$ satisfies to the unique extension property for derivations (Property (UDE)); see the corresponding definition in § 3.
@@
In the high-dimensional case, it is natural to use induction. Specifically, when an algebra $A$ has a chain of subalgebras $\mathbb{C}=R_{0}\subset R_{1}\subset\cdots\subset R_{n}=A$ such that an iterative application of Theorem 1.1 is possible, we have a tool for proving that $A\to\widehat{A}$ is a homological epimorphism; see the introduction to [26]. This plan works well in some situations but unfortunately cannot be directly applied to all high-dimensional solvable Lie algebras. Indeed, such an algebra admits an iterated semidirect sum decomposition,
@@
where $\mathfrak{f}_{1},\ldots,\mathfrak{f}_{n}$ are $1$-dimensional. Moreover, this decomposition actually induces an decomposition of the universal enveloping algebra into an iterated Ore extension; see, e.g., [21, pp. 33–34, 1.7.11(iv)]. But, in this case, $R_{n-1}\to(R_{n-1})_{\delta}$ is not usually an Arens–Michael envelope and, moreover, $(R_{n-1})_{\delta}$ cannot be represented as an analytic Ore extension. So we need modifications.
@@
Our strategy is to apply Theorem 1.1 using iterations. But the scheme proposed in [26] needs to be changed, as is clear from the previous discussion.
1. We use iterated analytic smash products instead of analytic Ore extensions; see Theorem 2.5. (An analytic Ore extension with trivial twisting is an analytic smash product but when the twisting is non-trivial this is not the case; thus most of parts of Theorem 9.12 in [26] are not covered by our approach.) Some preliminaries on Hopf $\mathop{\widehat{\otimes}}$-algebras, analytic smash products and Arens–Michael envelopes are contained in § 2.
2. We use completions of $\mathbb{C}[z]$ other than $\mathcal{O}(\mathbb{C})$, namely, the algebras $\mathfrak{A}_{s}$ defined in (2.7). The essential point is that $\mathcal{O}(\mathbb{C})\to\mathfrak{A}_{s}$ and some induced homomorphisms of analytic smash product satisfy Property (UDE). To prove theses facts we apply an auxiliary result on pushouts; see Theorem 3.5 and preparatory results in § 2.
3. We get an Arens–Michael envelope only on the final step of iteration. So we have to replace Theorem 1.2 by a more general assertion on homomorphisms satisfying Property (UDE); see Theorem 4.6. Combining it with a decomposition result essentially proved in [9] (Theorem 2.5), we finally deduce Theorem 4.3.
@@ 2. Hopf $\mathop{\widehat{\otimes}}$-algebras, analytic smash products and Arens–Michael envelopes
Consider the bifunctor $(-)\mathbin{\widehat{\otimes}}(-)$ of complete projective tensor product on the category of complete locally convex spaces and Hopf algebras in the corresponding symmetric monoidal category. We call them Hopf $\mathbin{\widehat{\otimes}}$-algebras (read ‘topological Hopf algebras’); see [20] or [25]. Also, we deal with $\mathbin{\widehat{\otimes}}$-algebras and $\mathbin{\widehat{\otimes}}$-(bi)modules, i.e., complete locally convex associative algebras and (bi)modules with jointly continuous multiplication. We assume that each algebra contains an identity and each module is unital. A $\mathbin{\widehat{\otimes}}$-(bi)module over a $\mathbin{\widehat{\otimes}}$-algebra $A$ is referred as an $A$$\mathbin{\widehat{\otimes}}$-(bi)module. We also assume that $\mathop{\widehat{\otimes}}$-algebra homomorphisms preserve identity and are continuous (as well as $\mathbin{\widehat{\otimes}}$-module morphisms).
@@
In the Hopf algebra theory, a notation introduced by Sweedler is widely used instead of the tensor notation in cases when the latter is not convenient. It was noted in [1, § 2.4] that this notation can be generalized to topological Hopf algebras. A version of the generalized Sweedler notation sufficient for our purposes is described in [9] and we briefly recall it here.
In the classical Sweedler notation, the comultiplication $\Delta$ on a Hopf algebra is written as
$\Delta(h)=\sum h_{(1)}\otimes h_{(2)},$ (2.1)
where an arbitrary representation of $\Delta(h)$ is taken.
@@
In the case of a Hopf $\mathop{\widehat{\otimes}}$-algebra, (2.1) is a short form of the above formula.
@@
In the generalized Sweedler notation, it takes the same form as in the classical:
@@
By the iterated limit theorem [18, Chapter 2, p. 69, Theorem 4], both iterated limits in (2.2) can be replaced by a simple limit of a net. In particular, this means that the iteration of the comultipication can be written as
@@
Also, the antipode axiom for a Hopf $\mathop{\widehat{\otimes}}$-algebra takes the standard form
@@
Let $H$ be a Hopf $\mathbin{\widehat{\otimes}}$-algebra and $A$ a $\mathbin{\widehat{\otimes}}$-algebra that is a left $H$$\mathbin{\widehat{\otimes}}$-module. Then $A\mathbin{\widehat{\otimes}}A$ and $\mathbb{C}$ are left $H$$\mathbin{\widehat{\otimes}}$-modules with respect to the actions given by
@@
For a $\mathbin{\widehat{\otimes}}$-algebra homomorphism $\psi\!:H\to B$ consider the adjoint action of $H$ on $B$:
@@
Let $H$ be a Hopf $\mathbin{\widehat{\otimes}}$-algebra and $A$ an $H$$\mathbin{\widehat{\otimes}}$-module algebra. The analytic smash product $A\mathop{\widehat{\#}}H$ is defined as a $\mathbin{\widehat{\otimes}}$-algebra endowed with $\mathbin{\widehat{\otimes}}$-algebra homomorphisms $i\!:A\to A\mathop{\widehat{\#}}H$ and $j\!:H\to A\mathop{\widehat{\#}}H$ such that the following conditions hold.
(A) $i$ is an $H$$\mathbin{\widehat{\otimes}}$-module algebra homomorphism with respect to the adjoint action associated with $j$.
@@
In the case when $H=\mathcal{O}(\mathbb{C})$ and the action of $H$ is induced by a derivation, the analytic smash product coincides with the corresponding analytic Ore extension [26, Remarks 4.4 and 4.2]. The same holds in the pure algebraic case, i.e., when the action of $\mathbb{C}[z]$ is induced by a derivation.
@@
$\mathfrak{A}_{s}\!:=\Bigl{\{}a=\sum_{n=0}^{\infty}a_{n}z^{n}\!:\|a\|_{r,s}\!:=\sum_{n=0}^{\infty}|a_{n}|\,\frac{r^{n}}{n!^{s}}<\infty\quad\forall r>0\Bigr{\}},$ (2.7)
where $z$ is treated as a formal variable, and endow $\mathfrak{A}_{s}$ with the topology determined by the seminorm family $(\|\cdot\|_{r,s};\,r>0)$. Denote also $\mathbb{C}[[z]]$, the algebra of all formal power series in $z$, by $\mathfrak{A}_{\infty}$.
@@
Here we assume that $\mathbb{C}[z]$ is endowed with the strongest locally convex topology and hence it is a $\mathop{\widehat{\otimes}}$-algebra with respect to this topology [26, Proposition 2.3]. In what follows we assume that all associative algebras of countable dimension are endowed with the strongest locally convex topology.